# Evaluating the Impact Of Spatial Features Of Mobility Data and Index Choice On Database Performance

### v1

@@ 5. Discussion & Future Work
While MobilityDB could be considered a more suitable system under test (SUT) due to its focus on mobility data, we chose PostGIS as it is more widely used, and our evaluation is limited to spatial queries. MobilityDB, despite its spatiotemporal capabilities, does not introduce novel spatial indexing strategies and offers no significant advantage in purely spatial scenarios without temporal aspects. Future work should consider extending the evaluation to spatiotemporal queries, in which case MobilityDB or similar databases would be more appropriate.
@@
Our benchmark includes diverse datasets and query types to ensure general applicability. Nevertheless, certain edge cases and access patterns may not be fully represented. While we normalize datasets by segment count to ensure fairness, this may unintentionally bias the results. We mitigate this by using multiple configurations and repeating experiments. Future work could expand the dataset range and normalization strategies to further reduce bias.
@@ 6. Conclusion
In this paper, we regarded the performance of a popular spatial database and how the choice of index, data format, and dataset characteristics impact the performance of a database, as these factors have been shown to impact the performance of a database in previous work but have not been considered together in detail.
For this, we designed novel approximation methods to determine the degree of skew and overlap in a dataset, which can be used to determine key data characteristics. We use these in combination with a self-designed benchmark to evaluate the performance of a popular spatial database, using both synthetic and real-world datasets. Our findings showed that especially data format and index choice can have a strong impact both on read and write performance, with data characteristics possibly influencing a possible speedup to be gained when relying on a specific data format. While we did not find a correlation for our ANN coefficient and performance, we believe that it provides a novel insight into trajectory distribution and may be useful for further research with other benchmark configurations. Future work should focus on extending our benchmark to further databases and indexing strategies to regard if data characteristics impact the performance of other databases and indexing strategies, while also including further datasets.

### v2

@@ 5. Discussion & Future Work
While MobilityDB could be considered a more suitable SUT due to its focus on mobility data, we chose PostGIS as it is more widely used, and our evaluation is limited to spatial queries. MobilityDB, despite its spatiotemporal capabilities, does not introduce novel spatial indexing strategies and offers no significant advantage in purely spatial scenarios without temporal aspects. Future work should consider extending the evaluation to spatiotemporal queries, in which case MobilityDB or similar databases would be more appropriate.
@@
Our benchmark includes diverse datasets and query types to ensure general applicability. Nevertheless, certain edge cases and access patterns may not be fully represented. While we equalize datasets by segment count to ensure fairness, this may unintentionally bias the results. We plan on evaluating the impact of using other equalization strategies, such as trajectory length or amount of trajectories in future work. We mitigate this by using multiple configurations and repeating experiments. Future work could expand the dataset range and normalization strategies to further reduce bias.
@@ 6. Conclusion
In this paper, we regarded the performance of a popular spatial database and how the choice of index, data format, and dataset characteristics impacted the performance of a database, as these factors have been shown to impact the performance of a database in previous work but have not been considered together in detail.
For this, we designed novel metrics and approximation methods to determine the degree of skew and overlap in a dataset, which can be used to determine key data characteristics. We use these in combination with a self-designed benchmark to evaluate the performance of a popular spatial database, using both synthetic and real-world datasets. Our findings showed that especially data format and index choice can have a strong impact both on read and write performance, with data characteristics possibly influencing a possible speedup to be gained when relying on a specific data format.
While we did not find a correlation for our ANN coefficient and performance, we believe that it provides a novel insight into trajectory distribution and may be useful for further research with other benchmark configurations. Future work should focus on extending our benchmark to further databases and indexing strategies to regard if data characteristics impact the performance of other databases and indexing strategies, while also including further datasets.